what is the equation of the line passing through (-2, 1) and having a slope of 1/2?

Answer by Alan3354(69443)

what is the equation of the line passing through (-2, 1) and having a slope of 1/2?
------------
Use
y-y1 = m*(x-x1) where (x1,y1) is the point and m is the slope.
y-1 = (1/2)*(x+2)
y = (1/2)x + 2 slope-intercept form
x - 2y = -4 standard form